OpenStudy (cenaida):

Which ordered pairs is the solution for the given system. x + 3y = 10 2x - y = 6 A. (4, 2) B. (7, 1) C. (16, -2) D. (3, 0)

OpenStudy (texaschic101):

do you prefer substitution or elimination ?

OpenStudy (texaschic101):

well...substitution is probably easier x + 3y = 10 x = -3y + 10 now sub -3y + 10 in for x in the other equation and solve for y. Once you have y, sub that back into either equation and find x.

OpenStudy (cenaida):

IS IT c?

OpenStudy (texaschic101):

let me check...hold on a sec

OpenStudy (texaschic101):

2x - y = 6 2(-3y + 10) - y = 6 -6y + 20 - y = 6 -7y + 20 = 6 -7y = 6 - 20 -7y = -14 y = 2 x + 3y = 10 x + 3(2) = 10 x + 6 = 10 x = 10 - 6 x = 4 No...answer is A (4,2)

OpenStudy (cenaida):

THANK YOU TEXASCHICK!

OpenStudy (texaschic101):

happy to help :)